OpenStudy (anonymous):

Can someone please explain how you would solve "log(7)343 as a logarithm with base 7"

OpenStudy (vishweshshrimali5):

See: \[\large{\log_{a} a^b = b}\]

OpenStudy (vishweshshrimali5):

Now, if you notice that: \[\large{343 = 7^3}\]

OpenStudy (vishweshshrimali5):

So, the problem you are asking for becomes: \[\large{\log_{7} 7^3}\] compare this with the formula: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 \[\large{\log_{a} a^b = b}\] \(\color{blue}{\text{End of Quote}}\) What do you think the answer must be ?

OpenStudy (anonymous):

there is also a nice method to cheat, if you do not know that \(7^3=343\)

OpenStudy (vishweshshrimali5):

Use a calculator ?

OpenStudy (vishweshshrimali5):

:D

OpenStudy (anonymous):

whip out mr calculator and type in \[\log(343)\div \log(7)\]

OpenStudy (vishweshshrimali5):

Hahaha :D Best method @satellite73 and probably the most used ^^^

OpenStudy (anonymous):

Oh..Now I get it. Thanks guys. I just didn't understand what numbers you would plug into the calculator

OpenStudy (vishweshshrimali5):

Great :)

OpenStudy (anonymous):

actually i was being silly it will work of course but this was not really a calculator exercise

OpenStudy (anonymous):

So what about if I had \[\log _{49}343\] with a base of 7

OpenStudy (anonymous):

if you want the calculator method, replace the \(7\) by \(49\) there is another method, but it requires some thinking

OpenStudy (anonymous):

I just need to know the calculator method for now but thank you. So you basically just ignore the "with a base of 7" part and use 49?

OpenStudy (vishweshshrimali5):

Actually what satellite is doing is the process called conversion of base.

OpenStudy (anonymous):

Yah isn't it the \[\log _{b}y\]=\[\frac{ \log _{y} }{ \log }\]

OpenStudy (vishweshshrimali5):

Well afraid not

OpenStudy (vishweshshrimali5):

It is basically this: \[\large{\log_a b = \cfrac{\log_c b}{\log_c a}}\]

OpenStudy (anonymous):

Oh. Thank you so much!